substitution method
-x+3y=10
2x+8y=-6
I need help with this problem if anyone could help, I would appreciate it.
Thanks!

Answer by ptaylor(2198)

substitution method
-x+3y=10----------------------eq 1
2x+8y=-6-------------------------eq2
solve for x in eq 1 and plug the value for x into eq 2:
-x+3y=10 subtract 3y from both sides
-x+3y-3y=10-3y
-x=10-3y multiply both sides by -1
x=3y-10 plug this into eq2
2(3y-10)+8y=-6 get rid of parens
6y-20+8y=-6 add 20 to both sides and collect like terms
14y-20+20=-6+20
14y=14 divide both sides by 14
y=1 substitute y=1 into eq 2-----------------------------ans for y
2x+8(1)=-6
2x+8=-6 subtract 8 from both sides
2x+8-8=-6-8 collect like terms
2x=-14 divide both sides by 2
x=-7---------------------------------------------ans for x
CK
-x+3y=10----------------------eq 1
7+3=10
10=10
2x+8y=-6-------------------------eq2
2*(-7)+8=-6
-14+8=-6
-6=-6


Hope this helps-----ptaylor